OpenStudy (anonymous):

what is 920 divided by 29510233720

OpenStudy (amistre64):

it is what it is :)

OpenStudy (anonymous):

what you mean

OpenStudy (amistre64):

i mean what i mean ...

OpenStudy (anonymous):

ok

myininaya (myininaya):

there is no way a teacher would ask you to divide this without a calculator

OpenStudy (anonymous):

its on my homework online question

OpenStudy (anonymous):

I have to do it without a calculator

myininaya (myininaya):

to what digit are you suppose to approximate the fraction to?

myininaya (myininaya):

where does it ask you to round?

OpenStudy (anonymous):

Its no decimal or rounding

myininaya (myininaya):

what?

myininaya (myininaya):

920 -------------=0.00000something 29510233720

OpenStudy (amistre64):

♫♫ you gotta know when to hold em ♫♫ know when to fold em ♫♫ no when to walk away ♫♫ know when to ruuuunnnn .... ♫♫♫

myininaya (myininaya):

what is the question then?

OpenStudy (anonymous):

The questions says " Suppose a department store operates 870 store in one chain and 50 in another chain. In 2007 the department Store had sales approximately $29510,233,720. What is the average of amount of sales made by each 920 stores"

OpenStudy (anonymous):

then it say divide Type a whole number for the answer

OpenStudy (amistre64):

then if anything divide the $$$ by 920

OpenStudy (anonymous):

nevermind I figured it out

OpenStudy (anonymous):

what did you get ?

OpenStudy (anonymous):

32076341

OpenStudy (anonymous):

you did it without a calculator ? you got it right :)

OpenStudy (anonymous):

Yeah without calculating.

OpenStudy (anonymous):

\[\frac{920 }{29510233720}=\frac{1}{32076341}=3.11756257*10^{-8} \]